OpenStudy (anonymous):

Logarithm and Functions Problem:

OpenStudy (anonymous):

If f(x) =\[\log_{a} x \] and \[h \neq0\], show that : \[\frac{ f(x+h)-f(x) }{ h }=\log_{a} (1+\frac{ h }{ x })^{\frac{ 1 }{ h }}\]

ganeshie8 (ganeshie8):

As a start, take the left hand side of you need to show and plug in the given function

OpenStudy (anonymous):

thats the part I'm having trouble with

OpenStudy (anonymous):

do i just add h to log_a * x?

OpenStudy (anonymous):

for the first one? f(x+h)

ganeshie8 (ganeshie8):

\(\log_a \) is function notation like \(\sin \), it neesd an argument to make sense : \(\log_a (x)\) is the logarithm of \(x\) relative to base \(a\)

OpenStudy (anonymous):

so what would be the result of f(x+h) after you plug in f(x)?

ganeshie8 (ganeshie8):

simply replace \(x\) by \(x+h\) : \(f(x) = \log_a(x)\) \(f(x+h) = \log_a(x+h)\)

OpenStudy (anonymous):

ok I've plugged in everything now how do i prove it

ganeshie8 (ganeshie8):

\[\begin{align} \frac{ f(x+h)-f(x) }{ h } &= \dfrac{\log_a(x+h)-\log_a(x)}{h}\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

what properties do you know about logarithms ?

OpenStudy (anonymous):

most of them: Change of base, the sum and difference etc.

ganeshie8 (ganeshie8):

very good :) time to use them.. use difference of logarithms property on numerator for now

OpenStudy (anonymous):

ok then what

ganeshie8 (ganeshie8):

step by step, first do that and the next step might become obvious :)

OpenStudy (anonymous):

oh ok i got it

OpenStudy (anonymous):

thanks!

ganeshie8 (ganeshie8):

\[\begin{align} \frac{ f(x+h)-f(x) }{ h } &= \dfrac{\log_a(x+h)-\log_a(x)}{h}\\~\\ &=\dfrac{\log_a\left(\dfrac{x+h}{x}\right)}{h}\\~\\ &=\dfrac{\log_a\left(1+\dfrac{h}{x}\right)}{h}\\~\\ &=\log_a\left(1+\dfrac{h}{x}\right)^{\frac{1}{h}}\\~\\ \end{align}\]

OpenStudy (anonymous):

Can you help me on this one too : Find the domain of \[f(x)=\log_7(\frac{ x+1 }{ x })\]

ganeshie8 (ganeshie8):

logarithm can only digest positive numbers

ganeshie8 (ganeshie8):

so the stuff inside parenthesis must always stay positive : \[\dfrac{x+1}{x} \gt 0 \] solve \(x\)

OpenStudy (jhannybean):

Oh, nevermind..., my question to you: I mentally thought of it as:\[\frac{1}{h}\cdot \log_a\left(1+\frac{h}{x}\right)\]which was the "k" you just mentioned, haha. Sorry for the redundancy!

ganeshie8 (ganeshie8):

Ahh yes yes thats a pretty interesting property if we think about why it has to be true

OpenStudy (jhannybean):

This loosely reminds me of the exponential growth function.